teh sum of three numbers is 35. the first number is three less than the second and the third is two more than four times the second. what are the three numbers?
Answer by checkley77(12569)

lET X-3, X, 4X+2 BE THE 3 NUMBERS.
X-3+X+4X+2=35
6X=35+3-2
6X=36
X=36/6
X=6 THE MIDDLE NUMBER.
6-3=3 THE FIRST NUMBER.
4*6+2=24+2=26 THE LARGEST NUMBER.
PROOF:
6+3+26=35
35=35